# Path planning with moving obstacles using stochastic optimal control

### v1

@@ III Geometric Symmetry
where $\mathcal{S}_{i}$ is a partition of +×+×$\mathbb{R}_{+}\times\mathbb{R}_{+}\times\left[0,\pi\right]$. More precisely, write $\mathcal{D}=\cup_{j=1}^{N_{d}-1}\mathcal{D}_{j}=[d_{1},d_{N_{d}}]$, where $\mathcal{D}_{j}=[d_{j},d_{j+1}]$, such that $\mathcal{D}_{j}\cap\mathcal{D}_{\bar{j}}=\emptyset$, for $j\neq\bar{j}$. Similarly, we define $\mathcal{E}=\cup_{l=1}^{N_{e}-1}\mathcal{E}_{l}$, and $\Theta=\cup_{m=1}^{N_{\theta}-1}\Theta_{m}$. Then we let $\mathcal{S}_{i}=\mathcal{D}_{j}\times\mathcal{E}_{l}\times\Theta_{m}$, where $j\in\left\{1,2,\cdots,N_{d}-1\right\}$, $l\in\left\{1,2,\cdots,N_{e}-1\right\}$, $m\in\left\{1,2,\cdots,N_{\theta}-1\right\}$, and $i=(j-1)(N_{e}-1)(N_{\theta}-1)+(l-1)(N_{\theta}-1)+m$. This results in $p=(N_{d}-1)(N_{e}-1)(N_{\theta}-1)$ partitions.
We also restrict ourselves to the case that $R_{u}\in\{0,\bar{R}_{u}\}$, and $R_{w}\in\{0,\bar{R}_{w}\}$ to have a tractable problem. Our previous results apply also to this case. We also discretize $\alpha$ and $\gamma$, and define $\bar{\mathcal{U}}=\left\{\bar{R}_{u}\begin{bmatrix}\cos\left(\alpha)\right)\\ \sin\left(\alpha)\right)\end{bmatrix}:\alpha\in\left\{0,\frac{\pi}{n_{1}},\frac{2\pi}{n_{1}},\cdots,\frac{(2n_{1}-1)\pi}{n_{1}}\right\}\right\}\cup\left\{\begin{bmatrix}0\\ 0\end{bmatrix}\right\}$, $\bar{\mathcal{W}}=\left\{\bar{R}_{w}\begin{bmatrix}\cos\left(\gamma)\right)\\ \sin\left(\gamma)\right)\end{bmatrix}:\gamma\in\left\{0,\frac{\pi}{n_{2}},\frac{2\pi}{n_{2}},\cdots,\frac{(2n_{2}-1)\pi}{n_{2}}\right\}\right\}\cup\left\{\begin{bmatrix}0\\ 0\end{bmatrix}\right\}$, where $n_{1}$ and $n_{2}$ are two natural numbers. We can rewrite the set $\bar{\mathcal{W}}$ as $\{w^{0},w^{1},\cdots,w^{2n_{2}}\}$, in which $w^{i}$ is $\begin{bmatrix}\cos\left(\frac{i\pi}{n_{2}}\right)\\ \sin\left(\frac{i\pi}{n_{2}}\right)\end{bmatrix}$, for $10\leq i\leq 2n_{2}-$ 1, and $w^{i}$ is $\begin{bmatrix}0\\ 0\end{bmatrix}$ for $i=2n_{2}$. This discretization results in that our previous results only apply approximately. We refer the reader to [11] for more details about fitted value iterations.
@@
where $d_{+}^{s}$, $e_{+}^{s}$, and $\theta_{+}^{s}$ are defined as in (9). We iterate until $\delta_{k}\triangleq||a_{k+1}-a_{k}||_{\infty}$ becomes sufficiently small.
Algorithm 1 summarizes the fitted value iterations.
Algorithm 1 Fitted Value Iteration
1:Input: Incremental Cost $\bar{f}$, Functions $F_{d}$, $F_{e}$ and $F_{\theta}$, Integer $n_{2}$, Control Set $\bar{\mathcal{U}}$, Set $\bar{\mathcal{W}}$, Tolerance $\varepsilon>0$, $N_{r}$ Different Sample Points $(d^{s},e^{s},\theta^{s})$
@@
4:while $||a_{k+1}-a_{k}||_{\infty}>\varepsilon$ do
@@
V Effect of Probability Distribution and Constraints
In the proposed method, a crucial assumption that needs to be satisfied is the conditions regarding the probability distribution of $w$. In this section, we relax this condition and also consider constraints on the states. As discussed in Section II, if $V$ satisfies the Ballman equation in (4), then minimizing the right-hand side of (4) gives the optimal policy for the specific probability distribution of $w$. We use the optimal value function $V$ that satisfies the Bellman equation in (4) and formulate the following problem:
$u_{k}^{\star}=\operatorname{argmin}_{u_{k}}\mathbb{E}\left[f(h_{k},r_{k})+V(h_{k+1},r_{k+1})\right]$ (16)
subject $\text{subject to}\quad\begin{cases}r_{k+1}=F_{r}\left(r_{k},u_{k}\right)\\ h_{k+1}=F_{h}\left(h_{k},w_{k}\right)\end{cases},\quad k\geq 0$
@@ V Effect of Probability Distribution and Constraints
where $\mathcal{X}$ is the set of all constraints on $x_{k}$. The expectation is with respect to $w$, and the probability distribution of $w$ is arbitrary. At each time step, we solve the optimization problem (16) in real-time to find the control signal $u_{k}^{\star}$, and then we apply this control signal. Had there been no constraints on $x_{k}$ and the probability distribution had satisfied our previous assumptions, then this online procedure would provide the optimal feedback. However, when we have constraints on the states and the probability distribution does not satisfy our previous assumptions, we only obtain a sub-optimal solution. This approach reduces the impact of the curse of dimensionality at the price of a suboptimal solution.
VI Numerical Results
In this section, we consider a numerical example and we use the proposed method described in Algorithm 1 to find the value function. We let $N_{d}=95$, $N_{e}=65$, $N_{\theta}=26$ , and $\{d_{1},d_{2},\cdots,d_{61}\}=\{0,0.05,0.1,\cdots,3\}$, $\{d_{62},d_{63},\cdots,d_{95}\}=\{3.6,4.4,5.2,\cdots,30\}$. Also we consider $\{e_{1},e_{2},\cdots,e_{31}\}=\{0,0.1,0.2,\cdots,3\}$, $\{e_{32},e_{33},\cdots,e_{65}\}=\{3.6,4.4,5.2,\cdots,30\}$. Moreover we let $\{\theta_{1},\theta_{2},\cdots,\theta_{26}\}=\{0,\frac{\pi}{25},\frac{2\pi}{25},\cdots,\pi\}$. This results in $p=150,400$ partitions. We consider $N_{r}=300,800$ sample points, i.e., two sample points per partition. Additionally, we let $\lambda=0.001$, and $R=1$. We let $h_{0}=\begin{bmatrix}2\\ 6\end{bmatrix}$, $r_{0}=\begin{bmatrix}4\\ 12\end{bmatrix}$, and $t=\begin{bmatrix}4\\ 3\end{bmatrix}$. We consider a probability distribution for $w$ such that its expected value drifts upward and to the right. The sets $\bar{\mathcal{U}}$ and $\bar{\mathcal{W}}$ are defined in Section IV, where $\bar{R}_{u}=\bar{R}_{w}=1$, and we let $n_{1}=n_{2}=16$. Moreover, we define the set $\mathcal{X}$ as follows
$\mathcal{X}=\left\{\begin{bmatrix}h^{x}\\ h^{y}\\ r^{x}\\ r^{y}\end{bmatrix}\in\mathbb{R}^{4}:0\leq r^{x},r^{y},h^{x},h^{y}\leq 20\right\}$
We consider $10000$ different realizations of $w$. We compare the results of the proposed method with receding horizon $A^{\star}$. Table I shows the cost, percentage of collisions, and execution time per stage for these methods, averaged over all different realizations of $w$. The proposed method has a lower cost than receding horizon $A^{\star}$. Additionally, the average computation time per stage and percentage of collisions for the proposed method are less than those for receding horizon $A^{\star}$. We plot one of the samples, and Fig. 4 shows the trajectory of the robot and the moving obstacle for the different methods. From this figure, the trajectory of the robot when using the proposed approach makes a detour to avoid the moving obstacle. Fig. 5 depicts the control signals for the different methods. In this figure, the first component of $u$, denoted by $u^{x}$ is shown in the top plot, and the second component of $u$, denoted by $u^{y}$ is shown in the bottom plot for these two cases. From this figure, the optimal control signal takes different angles, to avoid colliding with the moving obstacle. However, for almost all time instances, the control signal of the $A^{\star}$ approach remains constant, resulting in an almost straight line. Fig. 6 shows the distances between the robot and the target, and between the robot and the moving obstacle for the different methods at each time instance. From this figure, the robot reaches the target faster in the $A^{\star}$ approach than in our proposed method. However, the distance between the robot and the moving obstacle is larger in our proposed method compared to the $A^{\star}$ approach for most of the time instances.
TABLE I: Comparison of proposed method with receding horizon $A^{\star}$
Method Cost Computation Time Per Stage [sec] Percentage of Collision
Proposed Method $3.1168$ $0.0107$ $0.01$
Receding Horizon $A^{\star}$ $4.1683$ $0.2183$ $46.07$
@@ VI Numerical Results
Figure 5: Control Signal for different methods
@@
VII Conclusions
In this paper, we introduced a method to compute an optimal path for a robot in an environment with a stochastically moving obstacle. We reduced the dimension of the domain of the value function. In this way, we have less problem with the curse of dimensionality. When we compare our method with $A^{\star}$, our proposed method outperforms $A^{\star}$. More precisely, our method has lower cost, less computational time per stage, and fewer collisions.

### v2

@@ III Geometric Symmetry
where $\mathcal{S}_{i}$ is a partition of $\mathbb{R}_{+}\times\mathbb{R}_{+}\times\left[0,\pi\right]$. More precisely, write $\mathcal{D}=\cup_{j=1}^{N_{d}-1}\mathcal{D}_{j}=[d_{1},d_{N_{d}}]$, where $\mathcal{D}_{j}=[d_{j},d_{j+1}]$, such that $\mathcal{D}_{j}\cap\mathcal{D}_{\bar{j}}=\emptyset$, for $j\neq\bar{j}$. Similarly, we define $\mathcal{E}=\cup_{l=1}^{N_{e}-1}\mathcal{E}_{l}$, and $\Theta=\cup_{m=1}^{N_{\theta}-1}\Theta_{m}$. Then we let $\mathcal{S}_{i}=\mathcal{D}_{j}\times\mathcal{E}_{l}\times\Theta_{m}$, where $j\in\left\{1,2,\cdots,N_{d}-1\right\}$, $l\in\left\{1,2,\cdots,N_{e}-1\right\}$, $m\in\left\{1,2,\cdots,N_{\theta}-1\right\}$, and $i=(j-1)(N_{e}-1)(N_{\theta}-1)+(l-1)(N_{\theta}-1)+m$. This results in $p=(N_{d}-1)(N_{e}-1)(N_{\theta}-1)$ partitions. We refer the reader to [12] for more details about fitted value iteration.
@@
[…]
V Effect of Probability Distribution and Constraints
In the proposed method, a crucial assumption that must be satisfied is Assumption 2. In this section, we relax this condition and also consider constraints on the states. Notice that under this assumption, we can use symmetry reduction and reduce the dimension of the value function. However, this assumption imposes a restriction on the probability distribution of the motion of the obstacle. In realistic scenarios, such a restriction is not practical, and therefore the assumption must be relaxed to make the proposed approach applicable to real problems.
We use the optimal value function $V$ that satisfies the Bellman equation in (4) and formulate the following $N$-step lookahead rollout problem [13]:
$\underset{u_{l}\in\mathcal{U}}{\operatorname{minimize}}\quad\mathbb{E}\left[V(h_{k+N},r_{k+N})+\sum_{l=k}^{N+k-1}f(h_{k},r_{k})\right]$ (16)
[…]
where $\mathcal{X}$ is the set of all constraints on $x_{k}$. The expectation is taken with respect to $w$, and the probability distribution of $w$ does not necessarily satisfy Assumption 2. Note that for given $h$, $r$, and a static target position $t$, $V(h,r)=W(d,e,\theta)$, where $d,e$, and $\theta$ are defined in Definition 1.
At each time step, we solve the optimization problem in (16) in real time to obtain the optimal control sequence $u_{l}^{\star}$. The first control input of this sequence is then applied, and the remaining elements are discarded. Had there been no constraints on $x_{k}$ and the probability distribution had satisfied our previous assumptions, then this online procedure would provide the optimal feedback for (2). However, when state constraints are included or Assumption 2 is violated, the approach provides a suboptimal solution. This approach reduces the impact of the curse of dimensionality at the price of a suboptimal solution.
Solving the optimization problem (16) is computationally intensive. To reduce this computational burden, we apply the certainty equivalence (CE) principle, in which the stochastic disturbances $w_{k}$ are replaced by their expected values [13]. Using this approximation, we formulate the following problem
$\underset{u_{l}\in\mathcal{U}}{\operatorname{minimize}}\quad\left[V(h_{k+N},r_{k+N})+\sum_{l=k}^{N+k-1}f(h_{k},r_{k})\right]$ (17)
[…]
where $\bar{w}_{l}=\mathbb{E}\left[w_{l}\right]$.
Next, let $V_{N,\lambda}^{\star}(h_{0},r_{0},t)$ and $V_{N\text{(CE)},\lambda}^{\star}(h_{0},r_{0},t)$ denote the optimal values of the optimization problems in (16) and (17), respectively, for a fixed target position $t$, an initial condition $\begin{bmatrix}h_{0}&r_{0}\end{bmatrix}^{\mathsf{T}}$, and a given tuning parameter $\lambda$. Recall that the stage cost is a combination of two competing costs, weighted by the parameter $\lambda$ as in (3). To evaluate the effect of $\lambda$, we compute the expected performance over a representative set of initial states and targets. For each value of $\lambda$, we define
$\bar{V}_{N,\lambda}=\mathbb{E}\left[V^{*}_{N,\lambda}(h_{0},r_{0},t)\right]$ (18)
$\bar{V}_{N\text{(CE)},\lambda}=\mathbb{E}\left[V_{N\text{(CE)},\lambda}^{*}(h_{0},r_{0},t)\right]$
where the expectation is taken over a grid of sampled initial conditions and target positions. The purpose of this evaluation is to study the trade-off between the two performance criteria that form the cost in (3), namely the expected time to reach the target and the expected minimum distance maintained from the moving obstacle. By varying $\lambda$, we obtain the trade-off curve between the two performance criteria.
VI Numerical Results
In this section, we consider a numerical example and we use the proposed method described in Algorithm 1 to find the value function for different values of $\lambda$.
We let $N_{d}=115,\;N_{e}=85,\;N_{\theta}=26\!$, and $\{d_{1},d_{2},\cdots,d_{61}\}=\{0,0.05,0.1,\cdots,3\}\!$, $\{d_{62},d_{63},\cdots,d_{115}\}=\{3.5,4,4.5,\cdots,30\}\!$. Also we consider $\{e_{1},e_{2},\cdots,e_{31}\}=\{0,0.1,0.2,\cdots,3\}\!$, $\{e_{32},e_{33},\cdots,e_{85}\}=\{3.5,4,4.5,\cdots,30\}\!$. Moreover we let $\{\theta_{1},\theta_{2},\cdots,\theta_{26}\}=\{0,\frac{\pi}{25},\frac{2\pi}{25},\cdots,\pi\}\!$. This results in $p=239,400$ partitions. We consider $N_{r}=718,200$ sample points, i.e., three sample points per partition. Additionally, we let $R=1$, $\epsilon=10^{-8}$ , $\epsilon_{\text{tol}}=10^{-5}$, and $K=20$. The sets $\mathcal{U}$ and $\mathcal{W}$ are defined in Section II, and we let $n_{1}=n_{2}=16$. Moreover, we define the set $\mathcal{X}$ as follows
$\mathcal{X}=\left\{\begin{bmatrix}h^{x}&h^{y}&r^{x}&r^{y}\end{bmatrix}^{T}\in\mathbb{R}^{4}:0\leq r^{x},r^{y},h^{x},h^{y}\leq 20\right\}$
We implement Algorithm 1 in MATLAB and run it over a set of weighting parameters $\lambda\in[0,1]$. We define a discrete probability distribution on $\mathcal{W}$ by assigning the unnormalized weights
$\tilde{p}^{j}=\begin{cases}100,\quad\text{if }w^{i_{x}}>0,\text{ and }w^{i_{y}}>0\\ 1,\quad\text{otherwise}\end{cases}$
and normalizing them as $P(w^{i})=\frac{\tilde{p}^{i}}{\sum_{j=1}^{2n_{2}+1}\tilde{p}^{j}}$.
Algorithm 1 Reduced-Space Fitted Value Iteration
1:Input: Incremental Cost $\bar{f}$, Functions $F_{d}$, $F_{e}$ and $F_{\theta}$, Integer $n_{2}$, Control Set $\mathcal{U}$, Set $\mathcal{W}$, Tolerance $\epsilon_{tol}>0$, Maximum Number of Iterations $K$, $N_{r}$ Different Sample Points $(d^{s},e^{s},\theta^{s})$
@@ VI Numerical Results
4:while $a_{k+1}-a_{k}||_{\infty}>\epsilon_{\text{tol}} &$ $k<K$ do
@@
For the CBF approach [16], at each time instance, we solve the following expectation-based optimization problem
@@
$\text{subject to}\quad\mathbb{E}\left[B\left(F_{h}(h_{k},w_{k}),F_{r}(r_{k},u)\right)\mid x_{k}\right]\geq\alpha B(x_{k})$
[…]
$\text{subject to}\quad B\left(F_{h}(h_{k},\bar{w}_{k}),F_{r}(r_{k},u)\right)\geq\alpha B(x_{k})$
where $\bar{w}_{k}=\mathbb{E}[w_{k}]$. The nominal controller $u_{\text{nom}}$ is defined as the controller that drives the robot toward the goal as fast as possible without considering the obstacle.
We perform $50$ independent simulations, where in each trial the target position is sampled uniformly at random from the grid. The initial positions of the robot and the moving obstacle are also sampled uniformly at random, subject to the constraints $r_{0}-t||_{2}>1$ and $h_{0}-r_{0}||_{2}>1$. For each set of initial conditions, we generate $10$ independent realizations of the stochastic disturbance $w$ to approximate the costs in (16) and (17), i.e., $V_{N,\lambda}^{\star}(h_{0},r_{0},t)$ and $V_{N\text{(CE)},\lambda}^{\star}(h_{0},r_{0},t)$. For the CBF method, we evaluate the performance for multiple values of $\alpha\in(0,1)$, and $d_{0}$. Fig. 3 shows the trade-off curve obtained by varying the weighting parameter $\lambda$ compared with the receding horizon $A^{*}$, CBF, and CBF (CE) methods. Recall that we maximize the expected minimum distance between the robot and the obstacle. To obtain a standard Pareto trade-off plot, where both axes represent minimization, we display its negative value in Fig. 3. Thus, points closer to the origin correspond to better performance in both objectives. From Fig. 3, we observe that the best trade-off between the competing costs occurs at $\lambda^{*}=5\times 10^{-6}$. Additionally, from this figure, the proposed method achieves a lower trade-off curve on average over the grid, compared to CBF and CBF (CE). This means that, for different horizon lengths, the rollout solutions lie in the lower-left region of the plot relative to the curves of CBF and CBF (CE), indicating better performance for both objectives. Moreover, Fig. 3 shows that increasing the horizon shifts the trade-off curve further down and to the left, meaning that the longer horizon yields a better trade-off between the objectives (lower values for both costs) compared to curves that lie above and to the right. Furthermore, from this figure, the receding horizon $A^{*}$ approach always yields a solution that prioritizes minimizing the expected time to reach the target, as its points lie toward the lower-right part of the plot.
Figure 3: Trade-off curve between expected time to target (horizontal axis) and negative minimum distance to the obstacle (vertical axis) for different horizons, compared with receding horizon $A^{*}$, CBF, and CBF (CE) over the grid
Table I summarizes the expected time to reach the target and the expected minimum distance to the target for favorable values of $\gamma$, $\alpha$, and $d_{0}$ obtained from Fig. 3. From this table, the receding horizon $A^{*}$ method provides the fastest time to reach the target, while also achieving the smallest minimum distance among all methods. In contrast, the CBF and CBF (CE) methods exhibit a larger expected time to reach the target. Using the proposed method with different horizon lengths results in a balance between reaching the target efficiently and maintaining a larger minimum expected distance to the target, illustrating the trade-off between performance and safety.
Now, we let $h_{0}=\begin{bmatrix}2&6\end{bmatrix}^{T}$, $r_{0}=\begin{bmatrix}4&12\end{bmatrix}^{T}$, and $t=\begin{bmatrix}4&3\end{bmatrix}^{T}$. We consider $100$ different realizations of $w$, and compare the proposed method with the receding-horizon $A^{\star}$, CBF, and CBF (CE) approaches. Fig. 4 illustrates the trade-off curve obtained by varying the weighting parameter $\lambda$ compared with the other methods for the given initial condition and for different combinations of $\alpha\in(0,1)$, and $d_{0}$. From Fig. 4, the best trade-off between the competing costs occurs at $\lambda^{*}=5\times 10^{-6}$, which is the same value that we obtained from Fig. 3. Note that Fig. 3 shows the trade-off curve computed over the entire grid, whereas Fig. 4 illustrates the trade-off curve corresponding to the given initial conditions. From Fig. 4, for the given initial condition, the proposed method achieves a lower trade-off curve than CBF and CBF (CE) for some values of $\lambda$, while for other values of $\lambda$ the curve lies above those of CBF and CBF (CE). This indicates that, the proposed method does not always outperforms the other methods. However, when averaged over the entire grid of initial conditions, as shown in Fig. 3, our method outperforms the receding horizon $A^{*}$, CBF, and CBF (CE) methods.
Figure 4: Trade-off curve between expected time to target (horizontal axis) and negative minimum distance to the obstacle (vertical axis) for different horizons, for different horizons compared with receding-horizon $A^{*}$, CBF, and CBF (CE) for the given initial conditions
From Fig. 4, we select $\lambda=\lambda^{\ast}=5\times 10^{-6}$ for our method, and $\alpha=0.75$ and $d_{0}=1$ for the CBF and CBF (CE) methods, as these values provide the best trade-off for the given initial condition. We then plot the trajectory and control signal for a sample run of the different methods. Fig. 5 shows the trajectory of the robot and the moving obstacle for the different methods. From this figure, the trajectory of the robot when using the proposed approach makes a detour to avoid the moving obstacle. The trajectories produced by the CBF and CBF (CE) methods are identical in this example, and they also take a detour in order to satisfy the barrier function constraint. In contrast, the receding-horizon $A^{*}$ approach does not account for the moving obstacle and moves directly toward the goal without considering it. Fig. 6 shows the distance between the robot and the target, and the distance between the robot and the moving obstacle, over time for the different methods. From this figure, we notice that the robot reaches the target faster under the $A^{*}$ approach than under our proposed method, CBF, or CBF (CE). However, the distance between the robot and the moving obstacle is generally larger when using our proposed method compared to the $A^{*}$ approach, and is comparable to the distances achieved by the CBF and CBF (CE) methods. These results indicate that our method performs better than the other methods.
@@
VII Conclusions
In this paper, we used geometric symmetry to solve path planning problem for a robot in an environment with a stochastically moving obstacle. Compared with receding-horizon $A^{*}$, CBF, and CBF (CE), the proposed method achieves a better trade-off between reaching the target and maintaining safety; specifically, the trade-off curves lie below those of the other methods. Future work includes extending this framework to three-dimensional environments or to scenarios with moving targets. In the three-dimensional case with a static target, the symmetry reduction decreases the state dimension from $6$ to $3$. Additional directions include applying the reduced dimension value function within a reinforcement learning framework to learn the optimal value function in unknown environments, and generalizing the approach to multiple moving obstacles.
TABLE I: Comparison of proposed method with receding horizon $A^{\star}$, CBF, and CBF (CE), for $\lambda=5\times 10^{-6}$, $\alpha=0.75$, $d_{0}=1$
Method Expected Time to Reach the Target Expected Minimum Distance to Obstacle
Rollout ($N=1$) $10.278$ $8.0034$
Rollout ($N=1$, CE) $10.67$ $8.0377$
Rollout ($N=2$, CE) $11.388$ $8.2199$
Rollout ($N=3$, CE) $10.804$ $8.2337$
CBF $29.824$ $7.2857$
CBF (CE) $29.14$ $7.3081$
Receding Horizon $A^{\star}$ $8.896$ $7.3332$